OpenStudy (anonymous):

Please Help! Simplify the radical expression. 2sqrt*7+5sqrt*28.

OpenStudy (phi):

\[ 2 \sqrt{7} +5 \sqrt{28} \] you should factor the 28, and look for *pairs* of numbers remember that \( \sqrt{x\cdot x} \) simplifies to x example: \( \sqrt{4} = \sqrt{2\cdot 2}\) and we can simplify the square root of 2*2 to just 2 \( \sqrt{4} = \sqrt{2\cdot 2} =2\) example 2: \[ \sqrt{12} = \sqrt{2\cdot 2 \cdot 3} = \sqrt{2\cdot 2} \sqrt{3}= 2 \sqrt{3}\]

OpenStudy (phi):

can you find the factors of 28 ? are there any pairs ?

OpenStudy (anonymous):

4 and 7

OpenStudy (anonymous):

I actually got the answer 11sqrt7 for that problem. I asked for help on the wrong problem.. the problem I need help with is sqrt50-sqrt18-3sqrt12.

OpenStudy (solomonzelman):

\[2\sqrt{7}+5\sqrt{28}\]\[2\sqrt{7}+5\sqrt{7 \times 4}\]\[2\sqrt{7}+(5\sqrt{7} \times \sqrt{4})\]\[2\sqrt{7}+(5\sqrt{7} \times 2)\]\[2\sqrt{7}+10\sqrt{7}= 12\sqrt {7} \]

OpenStudy (anonymous):

dangit! I had it wrong?! lol Thank you!

OpenStudy (anonymous):

Ok. So I have 50sqrt2. 9sqrt2? Is this right?

OpenStudy (solomonzelman):

\[\sqrt{50}-\sqrt{18}-3\sqrt{12}\]\[5\sqrt{2}-3\sqrt{2}-(3 \times 2)\sqrt{2}\]\[5\sqrt{2}-3\sqrt{2}-6\sqrt{2}\]\[(5-3-6)\sqrt{2}\]\[-4\sqrt{2}\]

OpenStudy (solomonzelman):

see how I am taking the perfect squares out of the square roots?

OpenStudy (solomonzelman):

what is your actual question, I keep misunderstanding you, sorry.

OpenStudy (anonymous):

The answer in the book is 2sqrt2-6sqrt3.. I understand how you are taking the perfect squares out though. Simplify the radical expressions. sqrt*50-sqrt*18-3sqrt*12

OpenStudy (solomonzelman):

I did your question above then, but the book is wrong.

OpenStudy (solomonzelman):

Either I misread the question, or you mis-wrote it, or the book is stupid.

OpenStudy (phi):

*** the problem I need help with is sqrt50-sqrt18-3sqrt12. Ok. So I have 50sqrt2. 9sqrt2? Is this right? *** \[ \sqrt{50} - \sqrt{18} - 3 \sqrt{12} \] each square root, one at at time: 50 factors into 2*25 = 2*5*5 <-- notice the pair of 5's. 18 = 2*9= 2*3*3 <-- Look! a pair of 3's 12= 2*6= 2*2*3 <--- a pair of twos. We can take a pair out of the square root, toss one of the pair, and put the remaining number "out front" For example: \( \sqrt{50} = \sqrt{2\cdot 5 \cdot 5} = 5 \sqrt{2} \) do that to the other square roots: \[ \sqrt{18} = \sqrt{2 \cdot 3 \cdot 3} = 3 \sqrt{2} \\ \sqrt{12} = \sqrt{2 \cdot 2 \cdot 3} =2 \sqrt{3} \] the original problem is now \[ \sqrt{50} - \sqrt{18} - 3 \sqrt{12} \\ 5 \sqrt{2}-3 \sqrt{2} +-3 \cdot 2 \sqrt{3} \] notice that you have 5 square roots of 2 take away 3 square roots of 2, leaving 2 square roots of 2 or \( 2 \sqrt{2} \) also -3*2 = -6 in front of the sqrt of 3. The final answer is \[ 2 \sqrt{2} -6 \sqrt{3} \]